A flat (unbanked) curve on a highway has a radius of 220.0 m. A car rounds the curve at a speed of 25 m/s. (a) What is the minimum coefficient of friction that will prevent sliding?



1
Expert's answer
2021-11-30T17:44:11-0500

The minimum force of friction will help us find the minimum coefficient of friction:


"f=mv^2/R,\\\space\\
\mu mg=mv^2/R,\\\space\\
\mu=v^2/(2rg)=0.14."